OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1120. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many two-step equations auditioned today? 1300 1120 708 412 Uh...LOL

hero (hero):

Hint: x = one step y = two step x + y = 1120 3x - 2y = 1300

OpenStudy (anonymous):

im not sure where to start.

hero (hero):

2nd Hint: Multiply the first equation by 2

hero (hero):

Then add both equations together, then solve for x

hero (hero):

Then subtract x from 1120 to find y

OpenStudy (anonymous):

so 412..?

OpenStudy (anonymous):

i dont know 708

hero (hero):

You're guessing which means you didn't even attempt to solve it on your own.

OpenStudy (anonymous):

well your explaining isnt very well.

hero (hero):

I'll prove to you that my explanations were indeed well. Watch this:

hero (hero):

Original Equations x + y = 1120 3x - 2y = 1300 Multiply the first equation by 2: 2x + 2y = 2240 Add both equations together: 2x + 2y = 2240 3x - 2y = 1300 So adding results in 5x = 3540 Solve for x: x = 708 Subtract that from 1120: 412 So the answer is 412